# Visibly Fair Mechanisms

### v1

@@ 4 Incentives
Note that our richness assumption on the message space $M_{i}$ ensures the existence of truthful message for every officer. That is, for every preference $\succsim_{i}\in\mathcal{Q}_{i}$, there exists at least one message $m_{i}\in M_{i}$ that is truthful.
@@
We next define two conditions that are sufficient for a visibly fair mechanism to be strategy-proof, and weakening one of the two conditions are also necessary for a visibly fair mechanism to be strategy-proof.
The first condition ensures that an officer can provide preference information about other states he can obtain by deviating from a truthful message.
@@
Let $\psi$ be a mechanism and let $m$ be a message profile. The interest into state $s\in S\setminus\{\psi(m)\}$ is expressed by officer $i$ under $\psi(m)$ if either $s\succ_{m_{i}}\psi(m)_{i}$ or $\psi(m)_{i}\succ_{m_{i}}s$.
A mechanism $\psi$ satisfies expressiveness if for all $i\in I$, $\succsim_{i}\in\mathcal{Q}_{i}$, for any $m_{i}\in M_{i}$ that is a truthful message for $\succsim_{i}$, and any other message $\hat{m}_{i}\in{M}_{i}$ s.t. $\psi(\hat{m}_{i},m_{-i})_{i}\neq_{i}\psi(m)_{i}$, we have that interest into state $\psi(\hat{m}_{i},m_{-i})_{i}$ was expressed by officer $i$ under $\psi(m)$.
The second condition captures the idea that an officer cannot manipulate the availability of states by deviating from a truthful message.
@@
Let $\psi$ be a mechanism and let $m$ be a message profile. A state $s\in S\setminus\{\psi(m)\}$ is available to officer $i$ under $\psi(m)$ if
|$\left|\{\,j\in I:\psi(m)_{j}=s\}\,\cap\,\{\,j:\pi(j)<\pi(i)\}\right|\;<\;q_{s}.$
A mechanism $\psi$ satisfies availability if for all $i\in I$, $\succsim_{i}\in\mathcal{Q}_{i}$, for any $m_{i}\in M_{i}$ that is a truthful message for $\succsim_{i}$, and any other message $\hat{m}_{i}\in{M}_{i}$ s.t. $\psi(\hat{m}_{i},m_{-i})_{i}\neq_{i}\psi(m)_{i}$, we have that state $\psi(\hat{m}_{i},m_{-i})_{i}$ was available to officer $i$ under $\psi(m)$.
Interestingly, availability is too strong of a condition for ensuring strategy-proofness. Indeed, while availability ensures that an officer cannot manipulate the availability of states by deviating from a truthful message, we only require that an officer cannot manipulate the availability of strictly preferred states relative to his assignment — this condition, together with expressiveness in necessary for a visibly fair mechanism to be strategy-proof.
@@
A mechanism $\psi$ satisfies weak availability if for all $i\in I$, $\succsim_{i}\in\mathcal{Q}_{i}$, for any $m_{i}\in M_{i}$ that is a truthful message for $\succsim_{i}$, and any other message $\hat{m}_{i}\in{M}_{i}$ s.t. $\psi(\hat{m}_{i},m_{-i})_{i}\succ_{i}\psi(m)_{i}$, we have that state $\psi(\hat{m}_{i},m_{-i})_{i}$ was available to officer $i$ under $\psi(m)$.
@@
To give some intuition behind the results we given an example of a strategy-proof mechanism violating availability — but satisfying weak availability and expressiveness.
@@
  • Officers $I=\{\,i_{1},i_{2}\}$ with priority $\pi(i_{1})<\pi(i_{2})$.
  • States $S=\{\,s_{1},s_{2}\}$, each with capacity $q_{s}=1$.
Consider a mechanism $\psi$ where $i_{1}$ can only submit a single message $m_{i_{1}}$ without any preference information and $i_{2}$ can either submit message $m_{i}:s_{1}-s_{2}$ or $m^{\prime}_{i}:s_{2}-s_{1}$. Finally, let $\psi(m_{i_{1}},m_{i_{2}})=(s_{2},s_{1})$ and $\psi(m_{i_{1}},m^{\prime}_{i_{2}})=(s_{1},s_{2})$.
Clearly, the mechanism is strategy-proof as $i_{1}$ cannot influence the outcome by submitting a different message, and $i_{2}$ always gets his top choice when submitting preferences truthfully. Moreover, the mechanism is visibly fair as $i_{1}$ gives no preference information and $i_{2}$ gets his top choice.
It is easy to verify that availability is violated as, e.g., suppose we have $\succsim_{i_{2}}:s_{1}-s_{2}$, then submitting message $m_{i}$ which is truthful for $\succsim_{i_{2}}$, we have another message $m^{\prime}_{i}$ that leads to a different outcome $\psi(m_{i_{1}},m^{\prime}_{i_{2}})_{i_{2}}=\{s_{2}\}\neq\psi(m_{i_{1}},m_{i_{2}})_{i_{2}}=\{s_{1}\}$ which is not available under $\psi(m_{i_{1}},m_{i_{2}})_{i_{2}}$.
Note that weak availability is not violated as under $s_{1}\succsim_{i_{2}}s_{2}$ we have $\psi(m_{i_{1}},m_{i_{2}})_{i_{2}}\succ_{i_{2}}\psi(m_{i_{1}},m^{\prime}_{i_{2}})_{i_{2}}$, and analogous under $s_{2}\succsim_{i_{2}}s_{1}$ — and truthful message $m^{\prime}_{i_{1}}$ — we have $\psi(m_{i_{1}},m^{\prime}_{i_{2}})_{i_{2}}\succsim_{i_{2}}\psi(m_{i_{1}},m_{i_{2}})_{i_{2}}$.
Overall this example illustrates how availability too strong a requirement for strategy-proofness.
Furthermore, we give a second example of a mechanism satisfying expressiveness but violating weak availability and thus strategy-proofness.
@@
  • Officers $I=\{\,i_{1},i_{2}\}$ with priority $\pi(i_{1})<\pi(i_{2})$.
  • States $S=\{\,s_{1},s_{2}\}$, each with capacity $q_{s}=1$.
Consider a mechanism $\psi$ where $i_{1}$ can only submit a single message $m_{i_{1}}$ without any preference information and $i_{2}$ can either submit message $m_{i}:s_{1}\succ_{m_{i}}s_{2}$ or $m^{\prime}_{i}:s_{2}\succ_{m^{\prime}_{i}}s_{1}$. Finally, let $\psi(m_{i_{1}},m_{i_{2}})=(s_{1},s_{2})$ and $\psi(m_{i_{1}},m^{\prime}_{i_{2}})=(s_{2},s_{1})$.
Clearly, the mechanism is not strategy-proof as $i_{2}$ always gets his second choice when submitting a truthful message and can get his first choice by simply reporting the opposite message. The mechanism is visibly fair as $i_{1}$ does not give any preference information and $i_{2}$ only reported preferred state is given to the higher priority officer $i_{1}$.
The mechanism also satisfies expressiveness as under any message $i_{2}$ gives full preference information, while $i_{1}$ has a single message automatically satisfying expressiveness. At the same time the mechanism violates weak availability as e.g. under $\succsim_{i_{2}}:s_{1}-s_{2}$ and truthful message $m_{i_{2}}$ we have another message $m^{\prime}_{i_{2}}$ s.t. $\psi(m^{\prime}_{i_{1}},m_{i_{2}})_{i_{2}}\succ_{i_{2}}\psi(m_{i_{1}},m^{\prime}_{i_{2}})_{i_{2}}$ and $\psi(m^{\prime}_{i_{1}},m_{i_{2}})_{i_{2}}$ not being available under $\psi(m_{i_{1}},m^{\prime}_{i_{2}})$.
Finally, an alternative condition that ensures that a mechanism is strategy proof is as follows:
@@
A mechanism satisfies consistency if for any message $m_{i}\in M_{i}$ and any other message $\hat{m}_{i}\in M_{i}$ such that $\psi(\hat{m}_{i},m_{-i})_{i}\neq\psi(m)_{i}$ we have $\psi(\hat{m}_{i},m_{-i})_{i}\not\in G(\{\psi(m)_{i}\}\cup\{\psi(\hat{m}_{i},m_{-i})_{i}\},m_{i})$.
@@
A mechanism is strategy-proof if and only if it satisfies consistency.
@@ 4.1 Two special message spaces revisited
In this section we show that, in its most general form, neither the class of partition priority mechanism nor the ranked partition priority mechanisms are strategy-proof. In Example 4.3 we show that in a zonal message space partition priority mechanisms can fail availability. Moreover, in Example 4.4 we show that in a zonal message space with rankings no visibly fair mechanism, thus including ranked partition priority mechanisms, might satisfy expressiveness.
@@
  • Officers $I=\{\,i_{1},i_{2},i_{3}\}$ with priority $\pi(i_{1})<\pi(i_{2})<\pi(i_{3})$.
  • States $S=\{\,s_{1},s_{2},s_{3}\}$, each with capacity $q_{s}=1$.
  • A zonal message space with zones $z_{1}=\{s_{1},s_{2}\}$ and $z_{2}=\{s_{3}\}$.
  • Suppose each officer’s true preference is $s_{1}\succ_{i}s_{2}\succ_{i}s_{3}$. For each $i$, let $m_{i}$ be the message that ranks $s_{1}\succ_{m_{i}}s_{2}$, and let $m_{i}^{\prime}$ reverse it: $s_{2}\succ_{m_{i}^{\prime}}s_{1}.$
Consider the following zone selection function:
@@
Consider the following successful manipulation for $i_{3}$: Under the truthful profile $m$, the allocation is $(s_{1},s_{2},s_{3})$; but when $i_{3}$ flips her internal ranking $(m_{i_{3}}\!\to\!m_{i_{3}}^{\prime})$, the mechanism assigns $(s_{1},s_{3},s_{2})$, giving $i_{3}$ a better match $s_{2}\succ_{i_{3}}s_{3}$.
@@
Similarly, expressiveness is violated as e.g. officer $i_{3}$ does not express any preference information regarding $s_{2}$ and $s_{3}$ given message $m_{i_{3}}$, which is a truthful message for $\succ_{i_{3}}$, even though $s_{2}$ is allocated to $i_{3}$ under message $m_{i_{3}}^{\prime}$ and $s_{3}$ is allocated to $i_{3}$ under message $m_{i_{3}}$.
Finally, we give an example that illustrates the failure of a ranked partition priority mechanism to be strategy-proof due to its violation of expressiveness.
@@
  • Officers $I=\{i_{1},i_{2},i_{3}\}$ with priority $\pi(i_{1})<\pi(i_{2})<\pi(i_{3})$.
  • States $S=\{s_{1},s_{2},s_{3}\}$, each with capacity $q_{s}=1$.
  • A zonal message space ranking over zones: $z_{1}=\{s_{1},s_{2}\}$, $z_{2}=\{s_{3}\}$. Every officer’s message must fully rank $s_{1}$ and $s_{2}$ and also specify whether $z_{1}\rhd_{m_{i}}z_{2}$ or $z_{2}\rhd_{m_{i}}z_{1}$.
True preferences for $i_{1}$: $s_{1}\succ_{i_{1}}s_{2}\succ_{i_{1}}s_{3}$. This yields a unique truthful message $m^{\rhd}_{i_{1}}:(s_{1}\succ s_{2}\rhd s_{3})$, [^10] A message $m^{\rhd}_{i_{1}}:s_{1}\succ_{i_{1}}s_{2}\rhd s_{3}$ is our abbreviation for the preference message with $s_{1}\succ_{m_{i_{1}}}s_{2}$ and $z_{1}\rhd_{m_{i_{1}}}z_{2}$. ensuring $i_{1}$ obtains $s_{1}$ in any visibly fair mechanism.
@@
    Next, consider $\succ_{i_{2}}:\,s_{2}\succ s_{3}\succ s_{1}$. Here, two truthful messages are possible. One message $m_{i_{2}}:(s_{3}\rhd s_{2}\succ s_{1})$ again assigns $s_{3}$ to $i_{2}$. Another message $m^{\rhd\prime\prime}_{i_{2}}:(s_{2}\succ s_{1}\rhd s_{3})$ assigns $s_{2}$ to $i_{2}$.
Here, expressiveness is violated as e.g. officer $i_{2}$ does not express any preference information regarding $s_{2}$ and $s_{3}$ given message $m_{i_{2}}$, which is a truthful message for $\succ_{i_{2}}:\,s_{2}\succ s_{3}\succ s_{1}$, even though $s_{2}$ is allocated to $i_{2}$ under message $m^{\rhd\prime\prime}_{i_{2}}$ and $s_{3}$ is allocated to $i_{2}$ under message $m_{i_{2}}$. On the other hand, weak availability does not pose a problem in the above example.
@@
Consider the zonal message space with and without rankings, then both the partitioned priority mechanism and ranked partition priority mechanism might fail to be strategy-proof.

### v2

@@ 4 Incentives
Note that our richness assumption on the message space $M_{i}$ ensures the existence of a truthful message for every officer. That is, for every preference $\succsim_{i}\in\mathcal{Q}_{i}$, there exists at least one message $m_{i}\in M_{i}$ that is truthful.
@@
Strategy-proofness implies, therefore, that if an officer has multiple truthful messages, then they cannot lead to different outcomes. We next define two conditions that are sufficient for a visibly fair mechanism to be strategy-proof. Weakening one of the two conditions is also necessary for a visibly fair mechanism to be strategy-proof.
The first condition, expressiveness, requires that whenever an officer changes her message and thereby obtains another assignment, this new assignment must be comparable to the officer’s originally assigned state under the original message. Note that, the condition allows for the new assignment to be identical to the original assignment and that a state is always comparable to itself.
@@
Let $\psi$ be a mechanism and let $m$ be a message profile. The interest into state $s\in S$ is expressed by officer $i$ under $m$ if $s$ and $\psi(m)_{i}$ are comparable under $m_{i}$.
A mechanism $\psi$ satisfies expressiveness if for all $i\in I$, any $m_{i}\in M_{i}$ and any other message $\hat{m}_{i}\in{M}_{i}$ interest into state $\psi(\hat{m}_{i},m_{-i})_{i}$ was expressed by officer $i$ under $m$.
The second condition, availability, requires that whenever an officer changes her message and thereby obtains another assignment, this new assignment must always correspond to a state that was already available to her under the original message. Note that, the condition allows for the new assignment to be identical to the original assignment in which case the state is trivially available under the original message.
@@
Let $\psi$ be a mechanism and let $m$ be a message profile. A state $s\in S$ is available to officer $i$ under $m$ if $\left|\{\,j\in I:\psi(m)_{j}=s\}\,\cap\,\{\,j\in I:\pi(j)<\pi(i)\}\right|\;<\;q_{s}.$
A mechanism $\psi$ satisfies availability if for all $i\in I$, any $m_{i}\in M_{i}$ and any other message $\hat{m}_{i}\in{M}_{i}$ , we have that state $\psi(\hat{m}_{i},m_{-i})_{i}$ is available to officer $i$ under $m$.
Interestingly, availability is too strong of a condition for ensuring strategy-proofness. Indeed, while availability ensures that an officer cannot manipulate the availability of states by submitting a different message, weak availability only requires that an officer cannot manipulate the availability of weakly preferred states, evaluated at the original message, relative to her assignment under the original message. This condition, together with expressiveness in necessary for a visibly fair mechanism to be strategy-proof.
@@
A mechanism $\psi$ satisfies weak availability if for all $i\in I$, $\succsim_{i}\in\mathcal{Q}_{i}$, for any $m_{i}\in M_{i}$ that is a truthful message for $\succsim_{i}$, and any other message $\hat{m}_{i}\in{M}_{i}$ s.t. $\psi(\hat{m}_{i},m_{-i})_{i}\succsim_{i}\psi(m)_{i}$, we have that state $\psi(\hat{m}_{i},m_{-i})_{i}$ was available to officer $i$ under $m$.
@@
To give some intuition behind the results we give two examples. The first one is a strategy-proof mechanism violating availability — but satisfying weak availability and expressiveness.
@@
Consider a problem with two officers $I=\{i_{1},i_{2}\}$. Without loss of generality, we let officer $i_{1}$ have higher priority than $i_{2}$, i.e., $\pi(i_{1})<\pi(i_{2})$. There are two states $S=\{s_{1},s_{2}\}$, each with capacity $q_{s}=1$. Consider a mechanism $\psi$ where $i_{1}$ can only submit a single message $m_{i_{1}}$ without any preference information and $i_{2}$ can either submit message $m_{i_{2}}:s_{1}\succ_{m_{i_{2}}}s_{2}$ or $m^{\prime}_{i_{2}}:s_{2}\succ_{m^{\prime}_{i_{2}}}s_{1}$. Finally, let $\psi(m_{i_{1}},m_{i_{2}})=(s_{2},s_{1})$ and $\psi(m_{i_{1}},m^{\prime}_{i_{2}})=(s_{1},s_{2})$.
Clearly, the mechanism is strategy-proof as $i_{1}$ cannot influence the outcome by submitting a different message, and $i_{2}$ always gets her top choice when submitting preferences truthfully. Moreover, the mechanism is visibly fair as $i_{1}$ gives no preference information and $i_{2}$ gets her top choice.
It is easy to see that availability is violated as, e.g., consider message $m_{i}$ and message $m^{\prime}_{i}$. Note that, $m^{\prime}_{i}$ leads to a different outcome $\psi(m_{i_{1}},m^{\prime}_{i_{2}})_{i_{2}}=\{s_{2}\}\neq\psi(m)_{i_{2}}=\{s_{1}\}$ which is not available under $m$.
Note that weak availability is not violated as under $m_{i}$ which is a truthful message for $s_{1}\succ_{i_{2}}s_{2}$ we have $\psi(m)_{i_{2}}\succ_{i_{2}}\psi(m_{i_{1}},m^{\prime}_{i_{2}})_{i_{2}}$, and analogous under $m^{\prime}_{i_{1}}$ which is a truthful message for $s_{2}\succ_{i_{2}}s_{1}$ we have $\psi(m_{i_{1}},m^{\prime}_{i_{2}})_{i_{2}}\succ_{i_{2}}\psi(m)_{i_{2}}$.
Overall this example illustrates how availability is too strong a requirement for strategy-proofness.
The second example shows a mechanism satisfying expressiveness but violating weak availability and thus strategy-proofness.
@@
Consider a problem with two officers $I=\{i_{1},i_{2}\}$, where, without loss of generality, officer $i_{1}$ has higher priority than $i_{2}$, i.e., $\pi(i_{1})<\pi(i_{2})$. There are two states $S=\{s_{1},s_{2}\}$, each with capacity $q_{s}=1$. Consider a mechanism $\psi$ where $i_{1}$ can only submit a single message $m_{i_{1}}$ without any preference information and $i_{2}$ can either submit message $m_{i_{2}}:s_{1}\succ_{m_{i_{2}}}s_{2}$ or $m^{\prime}_{i_{2}}:s_{2}\succ_{m^{\prime}_{i_{2}}}s_{1}$. Finally, let $\psi(m_{i_{1}},m_{i_{2}})=(s_{1},s_{2})$ and $\psi(m_{i_{1}},m^{\prime}_{i_{2}})=(s_{2},s_{1})$.
Clearly, the mechanism is not strategy-proof as $i_{2}$ always gets his second choice when submitting a truthful message and can get his first choice by simply reporting the opposite message. The mechanism is visibly fair as $i_{1}$ does not give any preference information, and $i_{2}$ only reports the preferred state, which is given to the higher priority officer $i_{1}$.
The mechanism also satisfies expressiveness as under any message $i_{2}$ gives full preference information, while $i_{1}$ has a single message automatically satisfying expressiveness. At the same time the mechanism violates weak availability as e.g. under $m_{i_{2}}$ which is a truthful message for $s_{1}\succ_{i_{2}}s_{2}$ and another message $m^{\prime}_{i_{2}}$, where $\psi(m^{\prime}_{i_{1}},m_{i_{2}})_{i_{2}}\succsim_{i_{2}}\psi(m)_{i_{2}}$ we have that $\psi(m^{\prime}_{i_{1}},m_{i_{2}})_{i_{2}}$ is not available under $m$.
Finally, an alternative condition that ensures a mechanism is strategy-proof, without requiring visible fairness, is defined next. Coherence requires that whenever an officer changes her message and thereby obtains a different assignment, this new assignment is not in the set of undominated states — consisting of both the new and original assignment — evaluated at the original message. Note that this condition only applies when the two assignments are distinct states.
@@
A mechanism satisfies coherence if for any message $m_{i}\in M_{i}$ and any other message $\hat{m}_{i}\in M_{i}$ such that $\psi(\hat{m}_{i},m_{-i})_{i}\neq\psi(m)_{i}$ we have $\psi(\hat{m}_{i},m_{-i})_{i}\not\in G(\{\psi(m)_{i}\}\cup\{\psi(\hat{m}_{i},m_{-i})_{i}\},m_{i})$.
@@
A mechanism is strategy-proof if and only if it satisfies coherence.
@@ 4.1 Two special message spaces revisited
In this section we show that, neither the class of partition priority mechanism nor the ranked partition priority mechanisms are always strategy-proof.
In Example 4.3 we show that mechanisms using zonal message spaces can fail weak availability and expressiveness, and therefore not be strategy-proof. Moreover, in Example 4.4 we show that zonal message spaces can be sufficient conditions for visibly fair mechanisms to fail expressiveness, and therefore strategy-proofness.
@@
Consider a problem with three officers $I=\{i_{1},i_{2},i_{3}\}$. There are three states $S=\{s_{1},s_{2},s_{3}\}$, each with capacity $q_{s}=1$. The zonal message space for all officers consists of two zones: $z_{1}=\{s_{1},s_{2}\}$ and $z_{2}=\{s_{3}\}$. Suppose each officer’s true preference is $s_{1}\succ_{i}s_{2}\succ_{i}s_{3}$. For each officer $i$, let $m_{i}$ be the message that ranks $s_{1}\succ_{m_{i}}s_{2}$, and let $m_{i}^{\prime}$ reverse it: $s_{2}\succ_{m_{i}^{\prime}}s_{1}$.
Consider a zone selection function $\mathcal{C}$ that includes the following:
@@
Notice that these conditions are consistent with zone selection function that induces a partitioned priority mechanism. Consider the following successful manipulation for $i_{3}$: Under truthful report $m$, the allocation is $(s_{1},s_{2},s_{3})$; but when $i_{3}$ flips her internal ranking $(m_{i_{3}}\!\to\!m_{i_{3}}^{\prime})$, the mechanism assigns $(s_{1},s_{3},s_{2})$, giving $i_{3}$ a preferred assignment as $s_{2}\succ_{i_{3}}s_{3}$.
@@
Similarly, expressiveness is violated as e.g. officer $i_{3}$ does not express any preference information regarding $s_{2}$ and $s_{3}$, even though $s_{2}$ is allocated to $i_{3}$ under message $m_{i_{3}}^{\prime}$ and $s_{3}$ is allocated to $i_{3}$ under message $m_{i_{3}}$.
Finally, we give an example that illustrates that a ranked partition priority mechanism might not be strategy-proof due to its violation of expressiveness.
@@
Consider a setting with three officers $I=\{i_{1},i_{2},i_{3}\}$. The set of states is $S=\{s_{1},s_{2},s_{3}\}$, each with a capacity of $q_{s}=1$. The message space is zonal for all officers and requires officers to rank two zones: $z_{1}=\{s_{1},s_{2}\}$ and $z_{2}=\{s_{3}\}$. Each officer’s message must provide a full ranking over the states within zone $z_{1}$, i.e., between $s_{1}$ and $s_{2}$, and additionally indicate whether $z_{1}\rhd_{m_{i}}z_{2}$ or $z_{2}\rhd_{m_{i}}z_{1}$.
True preferences for $i_{1}$: $s_{1}\succ_{i_{1}}s_{2}\succ_{i_{1}}s_{3}$. This yields a unique truthful message $m^{\rhd}_{i_{1}}:(s_{1}\succ s_{2}\rhd s_{3})$, [^14] A message $m^{\rhd}_{i_{1}}:s_{1}\succ s_{2}\rhd s_{3}$ is our abbreviation for the preference message with $s_{1}\succ_{m_{i_{1}}}s_{2}$ and $z_{1}\rhd_{m_{i_{1}}}z_{2}$. ensuring $i_{1}$ obtains $s_{1}$ in any visibly fair mechanism.
@@
    Next, consider $\succ_{i_{2}}:\,s_{2}\succ s_{3}\succ s_{1}$. Here, two truthful messages are possible. One message $m_{i_{2}}:(s_{3}\rhd s_{2}\succ s_{1})$ again, as we have just argued, assigns $s_{3}$ to $i_{2}$. Another message $m^{\rhd\prime\prime}_{i_{2}}:(s_{2}\succ s_{1}\rhd s_{3})$ must assign $s_{2}$ to $i_{2}$.
Here, expressiveness is violated as e.g. officer $i_{2}$ does not express any preference information regarding $s_{2}$ and $s_{3}$ given message $m_{i_{2}}$, even though $s_{2}$ is allocated to $i_{2}$ under message $m^{\rhd\prime\prime}_{i_{2}}$ and $s_{3}$ is allocated to $i_{2}$ under message $m_{i_{2}}$. On the other hand, weak availability does not pose a problem in the above example.
@@
Consider the zonal message space with and without rankings, then both the partitioned priority mechanism and ranked partition priority mechanism might not be strategy-proof.